# An Experimental Design Framework for Label-Efficient Supervised Finetuning of Large Language Models

### v1

@@ 4.2 Evaluation on Standard Benchmarks
When comparing different experimental design (i.e. prompt selection) strategies to random sampling in Tables 1,2, we see almost dominant improvements (around 1% to 2%) on Facility Location ($\gamma$ = 0.002) strategy except on the 20K BBH-CoT case (where it closely aligns with that of random sampling). Compare to 90K budget where we get MMLU = $47.76(\pm 0.57)$, BBH-CoT = $40.49(\pm 0.30)$, we save approximately 50% annotation cost with respect to both classification (MMLU) and generation (BBH) tasks. Our hyperparameter-free kernel also exhibits improvement over random sampling, although the magnitude of improvement is not as pronounced as observed with the tuned kernel. This disparity could be attributed to the saturation of the Facility Location function, wherein the greedy order fails to generate a diverse summary conducive to downstream applications. In practice, since one could only choose one set of examples to annotate, we provide details in how to choose the hyperparameter in Section 4.3 before annotation and demonstrate the robustness through an ablation study. In addition, our proposed Min Margin strategy also gains much larger improvements compared to others, including the commonly used Mean Margin score.
On the other hand, we observe that most uncertainty-based selection approaches did not improve over random sampling. In some cases, they did not even exhibit monotonic improvements with respect to the selection size. This phenomenon has a well known explanation – uncertainty-based methods often annotate similar/redundant examples, which could potentially hinder to generalization performance when fine-tuning the model. This also suggests that a convex combination of uncertainty score with diversity-based selection or a two-stage selection as done in (Wei et al., 2015) may alleviate this problem and possibly further improve upon the gains.
4.3 Hyperparameter Selection for Facility Location
@@
In this section, we describe how to choose hyperparameter $\gamma$ before annotation begins. For a submodular function to be able to generate good summaries upon maximization, it should not saturate. That is, the facility location objective (equation (5)) starts to increase minimally by increasing the budget $k$, which suggests the new elements no longer remain representative of the downstream task. Formally, let $S_{1}\subset S_{2}\subset\ldots\subset S_{k}$ denote the greedy solution till size $k$ and $F(S_{k})=\sum_{i\in X}\max_{j\in S_{k}}w_{ij}$ denote the facility location function evaluation for set $S_{k}$ (defined in equation (5)). The gain by adding the $k$-th element is defined as $F(S_{k}|S_{k-1})\triangleq F(S_{k})-F(S_{k-1})$. In the case of RBF kernel, kernel width $\gamma$ controls the degree of saturation. Higher values of $\gamma$ lead to decreasing gain by adding additional elements. On the other hand, very low values such as $10^{-4}$ will result in the kernel becoming close to the diagonal matrix, and therefore not being able to capture the interactions between data points.
Therefore, to determine the potential range of kernel width, in Figure 2 we first visualize the gains of the submodular function when selecting 45K examples, for $\gamma\in\{10^{-3},5\times 10^{-3},10^{-2},10^{-1},1,10\}$. The observed trend reveals that for higher $\gamma$, gains tend to attain a very small value (and continue to decrease linearly) even before 1K (for $\gamma=10$) and 10K (for $\gamma=1$) elements are selected. Although reducing $\gamma$ helps, gains continue to decrease sublinearly ($\gamma=0.1$ after 20K). Notably, gains exhibit relative stability for $\gamma\in\{10^{-3},5\times 10^{-3},10^{-2}\}$ until we reach the desired budget of 45K, suggesting that we can safely choose $\gamma$ in between these values.
@@
With the range of potential $\gamma$ determined, we run an ablation study on 45K budget with four different kernel widths within this range to demonstrate the robustness of any hyperparameter in this range. As detailed in Table 3, the performance is consistently better than random selection strategy across different $\gamma$, which suggests the Facility location methods are less subject to hyper-parameter changes.
@@ 5 Conclusion and Future Work
  • Our setting finetunes the LLM only on the annotated instructions; future work could improve the utilization of unlabeled samples as done in Zhang et al. (2023).
@@ Acknowledgements
JZ and RN would like to thank the support of AFOSR/AFRL grant FA9550-18-1-0166, the UW-Madison Data Science Institute and the NSF grant DMS-2023239. GB, AD, and JB would like to thank NSF Grant Nos. IIS-2106937 and IIS-2148367, and by NIH/NHGRI U01 award HG009395.

### v3

@@ 4.2 Evaluation on Standard Benchmarks
When comparing different experimental design (i.e. prompt selection) strategies to random sampling in Tables 1,2, we see almost dominant improvements (around 1% to 2%) on Facility Location ($\gamma$ = 0.002) strategy except on the 20K BBH-CoT case (where it closely aligns with that of random sampling). Compared to 90K budget where we get MMLU = $47.76(\pm 0.57)$, BBH-CoT = $40.49(\pm 0.30)$, we save approximately 50% annotation cost with respect to both classification (MMLU) and generation (BBH) tasks. The facility location function with the cosine kernel also exhibits improvement over random sampling, although the magnitude of improvement is not as pronounced as observed with the tuned kernel. This disparity could be attributed to the saturation of the Facility Location function, wherein the greedy order fails to generate a diverse summary conducive to downstream applications. In practice, since one could only choose one set of examples to annotate, we provide details in how to choose the hyperparameter in Section 4.4 before annotation and demonstrate the robustness through an ablation study. In addition, our proposed Min Margin strategy also gains much larger improvements compared to others, including the commonly used Mean Margin score.
On the other hand, we observe that most uncertainty-based selection approaches did not improve over random sampling. In some cases, they did not even exhibit monotonic improvements with the selection size. This phenomenon has a well-known explanation – uncertainty-based methods often annotate similar/redundant examples, which could potentially hinder generalization performance when fine-tuning the model.
Figure 2: Evaluation by GPT-4 turbo by comparing model trained on 45K prompts selected by various strategies with the model trained on 90K random prompts. We use the win rate weighted by the continuous preferences of the GPT-4 turbo model. Error bars are reported as the standard errors across prompts.
Ablation
We now consider a combination of uncertainty-based selection with diversity, in the following Equation 4.
$\begin{split}S=\operatorname*{argmax}_{\begin{subarray}{c}S^{\prime}\subset X\\ |S^{\prime}|=k\end{subarray}}&\sum_{i\in X}\operatorname*{max}_{\begin{subarray}{c}j\in S^{\prime}\end{subarray}}w_{ij}\\ +&\operatorname{log}\left(1+\sum_{x\in S^{\prime}}\widetilde{U}_{\text{min margin}}(x)\right)\end{split}$ (4)
BBH-CoT MMLU
FL+MinMargin 40.71($\pm$0.63) 46.26($\pm$0.36)
MinMargin 39.66($\pm$0.67) 45.31($\pm$0.14)
FL 41.30($\pm$0.60) 47.63($\pm$0.24)
Table 3: Big-Bench-Hard chain-of-thoughts (BBH-CoT) and Massive Multitask Language Understanding (MMLU) evaluation of mixture of diversity and uncertainty strategy at the budget of 45K, averaged over 3 random seeds with standard error shown as confidence interval. We observe that the performance of the mixture interpolated between the two methods.
The first part governs the diversity and the latter part governs the uncertainty, for which, we consider the min-margin as the uncertainty score. Note that the objective in Equation 4 is a submodular maximization and hence we can use the greedy algorithm. This is a corollary of the fact that modular functions ($\sum_{x\in S^{\prime}}\tilde{U}(x)$) when composed with monotone, non-decreasing concave function is submodular, and submodularity is closed under conic combination Bilmes and Bai (2017); Lin and Bilmes (2011); Bhatt et al. (2024). In general, one can apply any monotone non-decreasing concave function, which becomes a design choice. Here, we consider $\operatorname{log}(1+x)$ which was chosen by Das et al. (2023). We consider min-margin for the uncertainty and RBF kernel based facility location. Table 3 shows the BBH-CoT and MMLU metrics for the mixture of min-margin and diversity at a budget of 45K and observe the performance of the mixture interpolated between the two methods. We posit that with the right choice of uncertainty metric, it is possible to improve beyond solely using uncertainty and diversity when mixed appropriately.
@@
4.3 Evaluation by GPT-4
To further demonstrate the annotation cost savings or various experimental design methods, we evaluate models finetuned on 45K prompts against the model trained on 90K prompts using random sampling with GPT-4 turbo as the judge, which displays a high agreement rate with ground truth human annotations . We follow the AlpacaEval framework (Li et al., 2023b) and report the win rate against the model finetuned on 90K randomly selected prompts. The models are evaluated on the 805 prompts from the standard AlpacaEval set (Li et al., 2023b). As showin in Figure 2, multiple strategies reliably achieve the same or better performance while only taking 50% of the annotation budget as compared to the model trained on 90K random prompts. Compared to result from standard benchmark, we again observed that Facility Location ($\gamma$ = 0.002) strategy gains dominating improvement. On the other hand, margin-based methods also exhibited significant improvements, while other uncertainty-based methods performed similarly or even worse than random sampling. This suggests that, although uncertainty-based methods may still be helpful in aligning with human preference, selecting the right metric is crucial. Overall, compared to standard benchmark results, the evaluation by GPT-4 turbo highlights even greater potential for experimental design in instruction finetuning. However, further investigation into how different strategies relate to various metrics (i.e., the model’s ability to perform different tasks) is needed to better understand the differences between standard and AlpacaEval benchmark results.
4.4 Hyperparameter Selection for Facility Location
In this section, we describe how to choose hyperparameter $\gamma$ before annotation begins. For a submodular function to be able to generate good summaries upon maximization, it should not saturate. That is, the facility location objective (equation (3)) eventually increases only minimally by increasing the budget $k$, which suggests the new elements no longer remain representative of the downstream task. Formally, let $S_{1}\subset S_{2}\subset\ldots\subset S_{k}$ denote the greedy solution up to size $k$ and $F(S_{k})=\sum_{i\in X}\max_{j\in S_{k}}w_{ij}$ denote the facility location function evaluation for set $S_{k}$ (defined in equation (3)). The gain by adding the $k$-th element is defined as $F(S_{k}|S_{k-1})\triangleq F(S_{k})-F(S_{k-1})$. In the case of RBF kernel, kernel width $\gamma$ controls the degree of saturation. Higher values of $\gamma$ lead to decreasing gain by adding additional elements. On the other hand, very low values such as $10^{-4}$ will result in the kernel becoming close to the diagonal matrix, and therefore not being able to capture the interactions between data points.
@@ 4.4 Hyperparameter Selection for Facility Location
Therefore, to determine the potential range of kernel width, in Figure 3 we first visualize the gains of the submodular function when selecting 45K examples, for $\gamma\in\{10^{-3},5\times 10^{-3},10^{-2},10^{-1},1,10\}$. The observed trend reveals that for higher $\gamma$, gains tend to attain a very small value (and continue to decrease) even before 1K (for $\gamma=10$) and 10K (for $\gamma=1$) elements are selected. Although reducing $\gamma$ helps, gains continue to decrease sublinearly ($\gamma=0.1$ after 20K). Notably, gains exhibit relative stability for $\gamma\in\{10^{-3},5\times 10^{-3},10^{-2}\}$ until we reach the desired budget of 45K, suggesting that we can safely choose $\gamma$ in between these values.
With the range of potential $\gamma$ determined, we run an ablation study on 45K budget with four different kernel widths within this range to demonstrate the robustness of any hyperparameter in this range. As detailed in Table 4, the performance is consistently better than random selection strategy across different $\gamma$, which suggests the Facility location methods are less subject to hyperparameter changes once the appropriate range for $\gamma$ has been identified.
5 Computational Complexity
We primarily used A100 and A40 GPUs. We pre-compute the embeddings which takes 12 hours. For A100, each trial takes roughly 24 GPU hours including the evaluation for 90K budget, 15 GPU hours for 45K and 9 hours for 30K respectively. For A40, roughly 30 GPU hours for 90K budget.
@@ 6 Conclusion and Future Work
  • Our setting finetunes the LLM only on the annotated instructions; future work could improve the utilization of unlabeled samples.
Limitations
While we successfully demonstrate the efficacy of experimental design on SFT, there are a few limitations of this work that we discuss. Firstly, we focus our experiments on the 7B version of Llama-2 since it is the most common and requires the least compute. Secondly, we conduct our experiments on curated datasets where we assume accurate responses to prompts are available. However, we do not consider the effect of inaccurate/noisy responses which can be prevalent when procuring responses from crowd-workers. Finally, we evaluate our model after SFT, and do not consider the effect of reinforcement learning with human feedback Ouyang et al. (2022) which is frequently applied after SFT.
Ethics Statement
Our work proposes a computationally inexpensive method of reducing the number of labeled instructions needed for SFT while preserving the performance of the resulting LLM. In doing so, we also mitigate the carbon footprint associated with training such models, which is critical given growing concerns about the environmental impact of modern ML systems. Furthermore, researchers and organizations with limited compute and insufficient financial resources to procure a vast number of crowd-workers can benefit from our work.
It is important to note, however, that all of the experimental design techniques that we propose are forms of biased sampling. If the subset of data chosen for annotation is not representative, it is possible biases of pretrained model will be amplified upon finetuning. Thus, it is critical for practitioners to ensure that the subsets selected by our experimental design framework are inclusive.
@@ Acknowledgements
JZ and RN would like to thank the support of AFOSR/AFRL grant FA9550-18-1-0166, the UW-Madison Data Science Institute and the NSF grants DMS-2023239, CNS-2112471. GB, AD, and JB would like to thank NSF Grant Nos. IIS-2106937 and IIS-2148367, and by NIH/NHGRI U01 award HG009395.